A wire with mass 60.0 g is stretched so that its ends are tied down at points 96.0 cm apart. The wire vibrates in its fundamental mode with frequency 70.0 Hz and with an amplitude of 0.400 cm at the antinodes.

What is the speed of propagation of transverse waves in the wire?

Compute the tension in the wire.
Answer #1

(a) v = f*lamba

= 70*2*L

= 70*2*0.96

v = 134.4 m/s

(b) v = sqrt(F_tension / mu)

mu = m/l = 0.06/0.96 = 0.0625 kg/m

F_tension = (mu)(v)^2

= (0.0625)(134.4)^2

= 1128.96 N